Serum cholesterol is an important risk factor for coronary disease. Suppose that serum cholesterol is approximately normally distributed with mean 212.9 mg/dL and standard deviation 51.78 mg/dL. a. Cholesterol levels in the top 20% are classified as high enough to warrant risk for heart disease. What cholesterol level corresponds to this cut-off? -mg/dL cholesterol- b. Suppose cholesterol levels in the bottom 40% are classified as normal. What cholesterol level corresponds to a normal reading? o cholesterol- mg/dL C. Between what two values do the middle 48% of cholesterol readings fall? o between mg/dL and mg/dL Note: z-scores should include 2 decimal places; all other answers should be rounded to 3 decimal placeslabel and answer all parts
Answer #1

We have,

         μ = 212.9           σ 51.78

a) Consider,

              P(X >x) = 0.20

             r-21290.20 51.78 P(

              r-2123) _ 0.20 PZ51.78        ..............(1)

       From Z-table(standard normal table), P(Z>0.84)=0.20045=0.20     ..........(2)

            Compare (1) and (2), we get

                         r-2129 = 0.84 51.78

                         = 212.9 + 0.84(51.78)

                        T-256.395

   Hence,

                 Z = 0.84

              Cholesterol = 256.395 mg/dL

b) Consider,

             P(X < z) 0.40

             = 0.4 Pl くー51.78

              P(Z<rac{x-212.9}{51.78})=0.4

                P(Z>-(rac{x-212.9}{51.78}))=0.4             symmetry of Z ..............(3)

        From Z-table, P(Z>0.25) =0.40129=0.4           .................(4)

                Compare (3) and (4) , we get

                   -(rac{x-212.9}{51.78})=0.25

                    r = 212.9-0.25(51.78)

                    199.955

     Hence,

               Z = 0.25

             Cholesterol = 199.955 mg/dL

c) Consider,

                -r < X < x) = 0.48

                 P(rac{-x-212.9}{51.78}<rac{X-mu}{sigma}<rac{x-212.9}{51.78})=0.48

                 P(rac{-x-212.9}{51.78}<Z<rac{x-212.9}{51.78})=0.48

                 P(Z>rac{-x-212.9}{51.78})-P(Z>rac{x-212.9}{51.78})=0.48

                 1-P(Z>rac{x-212.9}{51.78})-P(Z>rac{x-212.9}{51.78})=0.48

                 1-2P(Z>rac{x-212.9}{51.78})=0.48

                   P(Z>rac{x-212.9}{51.78})=rac{1-0.48}{2}

                   r-212.90.26 PZ51.78    .......................(5)

     From Z-table, P(Z>0.64)=0.26109 = 0.26....................(6)

                  Compare (5) and (6) , we get

                         rac{x-212.9}{51.78}=0.64

                       x=212.9+0.64(51.78)

                      x=246.039       ..................(*)

        Due to symmetry of normal,

                       x=212.9-0.64(51.78)

                       x=179.761    .................(**)

Hence,

                    Z=pm 0.64

               Between : 179.761 mg/dL and 246.039 mg/dL